OpenStudy (anonymous):

volume of solids of revolution help- solid bounded by y=x^2, x=2, y=0 revolved about x=2

OpenStudy (anonymous):

\[V=\int\limits_{0}^{4}\pi(\sqrt{y})^{2}dy\]=\[\frac{ \pi }{ 2 }y ^{2}]_{0}^{4}\]=\[8\pi \]